Session: Feb 2004 (session 1)
Wavelength: 18 cm


To give feedback on this experiment, click here

Experiment Feedback

Medicina (G.Maccaferri at Fri Feb 20 08:35:21 2004) Minor failures : POIN
Lost first 3 scans for antenna problems.

Onsala (Michael Lindqvist at Fri Feb 20 08:49:30 2004) Minor failures : RECO
FS did not switch Mark5-bank automatically. Lost some time.

Torun (K.Borkowski at Fri Feb 20 09:05:39 2004) Minor failures :
Occasional instabilities of signal levels in BBC5. No other known problems.

Jodrell_Bank (Alastair Gunn at Mon Feb 23 12:49:18 2004) Success :
No known problems.

Effelsberg (A. Kraus at Tue Feb 24 09:48:25 2004) Minor failures : RECO
Problems with a bad diskpack and with the MK5 system itself led to the loss of data between UT0106-0202 and UT0416-end.

Westerbork (G. Kuper at Wed Feb 25 10:05:35 2004) Minor failures : RECO RFI
RFI 1634.49 MHz (VC01 and VC02) Auto bank-switch failed. Done by hand.

Noto (P. Cassaro at Wed Feb 25 20:36:01 2004) Minor failures : LATE
The tape change at 6:08 UT took place too late. The observation restarted at 6:32 UT with 100 feet of difference in tape footage. Since 6:53 UT the tape footage has been totally corrected.

Robledo (Z.P. for Esther Moll at Wed Mar 10 10:05:00 2004) Minor failures : RECO
1. Time period allocated for this observation in Robledo was 00:15Z to 08:10Z. 2. Formatter was in sync. 3. This was a continuous and 2-head recording activity. Low tape sensor actuated unexpectedly at 03:44:07Z (pass 2R). Tape was manually repositioned by operator at beginning of next pass (3F). DR#102312 was opened. Just one scan was not recorded. No other tape problems for the rest of the observation. 3. Antenna arrived on time. No sources lost. 4. Video converter 5 (center frequency 1666.49 MHz) was affected by the known 1660-1670 MHz interference during the whole observation. Other RFIs present in some VCs: VC#1: schedule line 470 (scan 41) VC#3: schedules lines 301, 887 and 956 (scans 26, 77, 83) 5. The "onsource" status is SLEWING for most of the observation. This status is not real, actually the antenna was properly TRACKING and no sources were lost. 6. The calibration signal (noise diode) was not configured due to lack of time during Pre-experiment configurations. Nominal system temperature at zenith is 30.8 K.